two sums of y=x
y=-3x-8
I know on a linear graph they meet at co-ordinates (-2,-2)
ive been asked to solve the two above sums usng eliminaton and subsitution method for simultanious linear equations I have spent all day trying to work it out ad just cant grasp it could you help please

Answer by nilan(17)

y=x -------(A)
y=-3x-8 -------(B)
(A)-(B)
y-y = x-(-3x-8)
0 = x+3x+8
0 = 4x+8
dividing both side by 4
0 = x+2
so x = -2
substituting this into equation (A)
which is y=x
y=-2